OpenStudy (anonymous):

Find the inverse of the function y = x2 – 12.

OpenStudy (akashdeepdeb):

y = x2 - 12 Inverses are symmetrical with respect to line y=x on xy plane. So replace y with x and x with y in the equation you are given. We get x = y^2 - 12 y^2 = x+12 y = \(\sqrt{x+12}\) That is the inverse. Understood? :)